OpenStudy (anonymous):

find the slope if it exists, of the line containing the pair of points -8,-15 and -19,-16

OpenStudy (anonymous):

m= (y1-y)/(x1-x)=(-15-(-16))/(8-(-19))

OpenStudy (anonymous):

y2 - y1 -16 - (-15) -16 + 15 -1 1 slope m = --------- = ------------ = -------- = ----- = ---- x2 - x1 -19 - (-8) -19 + 8 -11 11

OpenStudy (anonymous):

@inik all points r negative, i think u hv made a mistake...

OpenStudy (anonymous):

i didn't see _ in a front of 8... Harkirat is correct!

OpenStudy (anonymous):

thx inik :)

OpenStudy (anonymous):

Harkirat: Hit the "Good Answer" button.

OpenStudy (anonymous):

for whom???????